# UniDepthV2: Universal Monocular Metric Depth Estimation Made Simpler

### v1

@@ IV-B Comparison with The State of The Art
Method SUNRGBD HAMMER IBims-1 TUM-RGBD
$\mathrm{\delta_{1}}\uparrow$ $\mathrm{F_{A}}\uparrow$ $\mathrm{\rho_{A}}\uparrow$ $\mathrm{\delta_{1}}\uparrow$ $\mathrm{F_{A}}\uparrow$ $\mathrm{\rho_{A}}\uparrow$ $\mathrm{\delta_{1}}\uparrow$ $\mathrm{F_{A}}\uparrow$ $\mathrm{\rho_{A}}\uparrow$ $\mathrm{\delta_{1}}\uparrow$ $\mathrm{F_{A}}\uparrow$ $\mathrm{\rho_{A}}\uparrow$
Metric3D†‡ [14] $1.9$ $0.9$ $75.1$ $7.7$
Metric3Dv2†‡ [16] $81.2$ $\mathbf{65.3}$ $68.4$ $63.0$
ZoeDepth [33] $80.9$ $0.9$ $49.8$ $55.6$
UniDepth [17] $94.3$ $78.6$ $85.8$ $1.8$ $52.1$ $55.3$ $15.7$ $30.3$ $\mathbf{76.6}$ $72.3$ $54.8$ $86.8$
MASt3R [74] $80.1$ $71.5$ $\underline{92.0}$ $2.2$ $38.1$ $\mathbf{86.5}$ $61.0$ $55.7$ $76.0$ $52.4$ $44.1$ $\underline{93.7}$
DepthPro [37] $83.1$ $71.1$ $89.3$ $29.4$ $\underline{71.0}$ $69.1$ $82.3$ $62.8$ $75.9$ $56.9$ $48.1$ $\mathbf{96.5}$
UniDepthV2-Small $90.8$ $74.2$ $87.7$ $20.1$ $52.6$ $77.5$ $86.6$ $62.4$ $67.5$ $69.0$ $50.6$ $86.1$
UniDepthV2-Base $\underline{94.4}$ $\underline{79.9}$ $91.1$ $30.6$ $57.0$ $65.6$ $\underline{89.7}$ $\underline{68.5}$ $\underline{76.5}$ $\underline{77.5}$ $\underline{57.3}$ $89.4$
UniDepthV2-Large $\mathbf{96.4}$ $\mathbf{84.6}$ $\mathbf{93.4}$ $\underline{64.5}$ $\mathbf{74.9}$ $\underline{78.3}$ $\mathbf{94.5}$ $\mathbf{70.9}$ $74.1$ $\mathbf{90.5}$ $\mathbf{62.9}$ $89.6$
@@
TABLE II: Results for Outdoor Domains. All methods are tested in a zero-shot fashion. Missing values (-) indicate the model’s inability to produce the respective output. †: Requires ground-truth (GT) camera for 3D reconstruction. ‡: Requires GT camera for 2D depth map inference.
Method ETH3D Sintel DDAD NuScenes
$\mathrm{\delta_{1}}\uparrow$ $\mathrm{F_{A}}\uparrow$ $\mathrm{\rho_{A}}\uparrow$ $\mathrm{\delta_{1}}\uparrow$ $\mathrm{F_{A}}\uparrow$ $\mathrm{\rho_{A}}\uparrow$ $\mathrm{\delta_{1}}\uparrow$ $\mathrm{F_{A}}\uparrow$ $\mathrm{\rho_{A}}\uparrow$ $\mathrm{\delta_{1}}\uparrow$ $\mathrm{F_{A}}\uparrow$ $\mathrm{\rho_{A}}\uparrow$
Metric3D†‡ [14] $19.7$ $1.4$ $81.9$ $75.4$
Metric3Dv2†‡ [16] $\mathbf{90.0}$ $\mathbf{34.5}$ $\underline{87.6}$ $84.1$
ZoeDepth [33] $33.8$ $5.6$ $27.9$ $33.8$
UniDepth [17] $18.5$ $27.6$ $42.6$ $13.2$ $40.2$ $65.6$ $85.8$ $\underline{72.8}$ $\mathbf{98.1}$ $84.6$ $\underline{64.4}$ $\mathbf{97.7}$
MASt3R [74] $21.4$ $28.4$ $\underline{92.2}$ $17.2$ $41.5$ $72.2$ $4.3$ $22.1$ $74.6$ $2.7$ $13.6$ $78.3$
DepthPro [37] $39.7$ $41.2$ $77.4$ $26.2$ $49.7$ $75.2$ $29.9$ $42.1$ $83.0$ $56.6$ $46.5$ $79.1$
UniDepthV2-Small $64.6$ $44.3$ $78.4$ $14.6$ $37.1$ $73.5$ $83.3$ $68.5$ $94.7$ $82.1$ $59.7$ $96.2$
UniDepthV2-Base $75.4$ $\underline{53.5}$ $91.4$ $31.9$ $\mathbf{51.8}$ $\underline{75.9}$ $86.8$ $71.4$ $96.1$ $\underline{85.3}$ $63.6$ $96.6$
UniDepthV2-Large $\underline{85.2}$ $\mathbf{59.3}$ $\mathbf{92.6}$ $\underline{34.4}$ $\underline{51.4}$ $\mathbf{76.3}$ $\mathbf{88.2}$ $\mathbf{73.3}$ $\underline{96.7}$ $\mathbf{87.0}$ $\mathbf{66.7}$ $\underline{97.2}$
@@
Figure 4: Invariance to image shape. UniDepthV2 is trained with a variable input shape pipeline in addition to random resizing for each of the image pairs. The proposed training strategy improves the robustness in terms of predicted depth scale and accuracy ($\delta_{1}$) to the input image’s shape compared to two other state-of-the-art methods.
@@
We evaluate our method on eight zero-shot validation sets, covering both indoor and outdoor scenes, as shown in Table I and Table II, respectively. Our model performs better than or at least on par with all baselines, even outperforming methods that require ground-truth camera parameters at inference time, such as [14, 16]. Notably, UniDepthV2 excels in 3D estimation, as reflected in the $\mathrm{F_{A}}$ metric, where it achieves a consistent improvement ranging from 0.5% to 18.1% over the second-best method. Additionally, it outperforms UniDepth [17] in nearly all cases, except for the $\mathrm{\rho_{A}}$ metric on IBims-1, DDAD, and NuScenes. This demonstrates that our proposed version is a significant step forward in both performance and efficiency. However, the camera parameter estimation ($\mathrm{\rho_{A}}$) sees only marginal improvements, indicating that the limited diversity of training cameras remains a challenge that could be addressed with additional camera-only training, as suggested in [37]. Table III and Table IV show results for models fine-tuned on the NYU and KITTI training sets and evaluated on their respective validation splits, following standard protocols. Fine-tuning performance serves as an indicator of a model’s ability to specialize to specific downstream tasks and domains. UniDepthV2 effectively adapts to new domains and outperforms methods that were pre-trained on large, diverse datasets before fine-tuning on NYU or KITTI, such as [33, 16, 29], This is particularly evident in the outdoor setting (KITTI), as shown in Table IV. As detailed in Section III-E, our training strategy incorporates variable image aspect ratios and resolutions within the same distributed batch. Combined with camera conditioning and invariance learning, this approach enhances the model’s robustness to changes in input image shape. Figure 4 quantifies this effect: the y-axis represents normalized metric accuracy ($\mathrm{\delta}_{1}$ scaled by the method’s maximum value), while the x-axis varies the image shape. The normalization ensures a consistent scale across models. UniDepthV2 is almost invariant to image shape, demonstrating that it can effectively trade off resolution for speed without sacrificing accuracy, as clearly illustrated in Figure 4.
@@
TABLE III: Comparison on NYU validation set. All models are trained on NYU. The first 4 are trained only on NYU. The last 4 are fine-tuned on NYU.
@@
UniDepthV2 $\underline{98.8}$ $\mathbf{99.8}$ $\mathbf{100}$ $\mathbf{4.68}$ $\mathbf{0.180}$ $\mathbf{0.020}$
TABLE IV: Comparison on KITTI Eigen-split validation set. All models are trained on KITTI Eigen-split training and tested on the corresponding validation split. The first 4 are trained only on KITTI. The last 4 are fine-tuned on KITTI.
@@
UniDepthV2 $\mathbf{98.9}$ $\mathbf{99.8}$ $\underline{99.9}$ $\mathbf{3.73}$ $\mathbf{1.71}$ $\mathbf{0.061}$
Figure 5: Confidence invariance. The uncertainty output of UniDepthV2 represents the predicted error. The confidence is obtained as the inverse uncertainty and the output is evaluated by taking into account only the pixels with a confidence higher than the corresponding x-axis. Y-axis reports the normalized RMSE to have a consistent scale among different datasets, where normalization involves dividing the RMSE by the value with threshold 0, namely evaluating over all pixels.
IV-C Ablation Studies
The importance of each new component introduced in UniDepthV2 in Sec. III is evaluated by ablating the method in Tables V, VI, and VII. All ablations exploit the predicted camera representation, if not stated otherwise. Table V evaluates the impact of various architectural modifications compared to UniDepth [17], analyzing their effects on both performance and efficiency. Table VI assesses the importance of the proposed loss function (Sec. III-D) and examines the effect of applying the geometric invariance loss originally introduced in UniDepth [17] (Sec. III-C) in different spaces. The rationale behind our design choices is to maintain simplicity while maximizing effectiveness. Additionally, in Table VII we analyze the role of camera conditioning and report results for the original UniDepth under the same training and evaluation setup as our method for a direct comparison. The evaluation is based on four key metrics: $\mathrm{\delta}_{1}$, which measures metric depth accuracy; $\mathrm{SI}_{\log}$, which assesses scale-invariant scene geometry; $\mathrm{F_{A}}$, which captures the 3D estimation capability; and $\mathrm{\rho_{A}}$, which evaluates monocular camera parameter estimation. All reported metrics correspond to the aggregated zero-shot performance across datasets, as detailed in Sec. IV-A.
@@ IV-C Ablation Studies
TABLE V: Architectural ablations. The different architectural additions (“+”) and subtractions (“-”) from the original UniDepth [17] are reported. “- SHE + Sine”: camera encoding via Sine encoding instead of Spherical Harmonic Transform of the pinhole-based pencil of rays. “- Attention”: attention layers in the decoder are removed. “+ ResNet Blocks”: the attention layers in the decoder are substituted with simpler ResNet blocks. “+ Multi-resol.”: the decoder has lateral connections with the shallower encoder layer, rather than a simpler merging of all resolutions in the bottleneck.
@@
5 + Multi-resol. $54.5$ $16.3$ $56.0$ $77.9$ 25.0 34.2
TABLE VI: Loss ablations. $\mathcal{L}_{\mathrm{EG-SSI}}$ refers to either employing or not the proposed Edge-Guided Normalized loss; $\mathbf{O}_{\mathcal{L}_{\mathrm{con}}}$ indicates the output there the geometry consistency loss is applied to.
@@
TABLE VII: Model ablations. The “Model” column refers to architecture and training strategy employed. “V1” is the original UniDepth, while “V2” is the proposed UniDepthV2. “Cond” specifies whether the camera-prompting mechanism is present or not.
@@
Architecture.  Table V outlines the key modifications that transform the original UniDepth [17] architecture into UniDepthV2. The first major change is the removal of spherical harmonics (SH)-based encoding, which is computationally inefficient. Instead, we revert to standard Sine encoding (row 2). While the difference in performance is minimal in our setup, we hypothesize that the encoding’s impact diminishes as the model benefits from larger and more diverse training data across different cameras. Next, we eliminate the attention mechanism in row 3 due to its high computational cost. This removal results in a significant performance drop, e.g. -4.3% for $\mathrm{\delta}_{1}$, but yields a greater than 2x improvement in efficiency. In row 4, we replace the pure MLP-based decoder with ResNet blocks, introducing spatial $3\times 3$ convolutions. This modification enhances performance by leveraging local spatial structure while inducing a minimal impact on efficiency. Finally, row 5 integrates a multi-resolution feature fusion from the encoder to the decoder, following an FPN-style design. This final architecture significantly reduces computational cost while preserving overall performance: the final model (row 5) achieves similar performance to the original UniDepth (row 1) while requiring only one-third of the computation. $\mathcal{L}_{\mathrm{EG-SSI}}$ Loss.  The effectiveness of the proposed $\mathcal{L}_{\mathrm{EG-SSI}}$ loss, detailed in Sec. III-D, is evaluated in row 2 vs. row 3 of Table VI. Introducing this loss results in a 4.7% improvement in $\mathrm{\delta}_{1}$ and a 1.8% improvement in $\mathrm{F_{A}}$, demonstrating its contribution to both metric accuracy and 3D estimation. Interestingly, despite $\mathcal{L}_{\mathrm{EG-SSI}}$ not explicitly supervising camera parameter estimation, the $\mathrm{\rho_{A}}$ metric also shows improvement. This suggests that the loss contributes to a less noisy training process, leading to better feature representations in the encoder. A qualitative comparison of the impact of $\mathcal{L}_{\mathrm{EG-SSI}}$ is presented in Fig. 6. The difference between the third and fourth columns highlights the visual impact of the proposed loss, particularly in refining depth discontinuities. Additionally, the comparison between the second and third columns illustrates the combined effect of architectural changes and increased data diversity, showing improved reconstruction of finer details, such as body parts that were previously smoothed or missed. $\mathcal{L}_{\mathrm{con}}$ Output Space.  UniDepthV2 introduces multiple instances of camera-conditioned depth features $\mathbf{D}|\mathbf{E}$, corresponding to different decoder resolutions, as described in Sec. III-E. This contrasts with the original UniDepth [17], which relied on a single conditioning point. Given this architectural shift, we argue that deep conditioning may not be optimal. Features at different resolutions encode varying levels of abstraction, and enforcing deep conditioning introduces additional design freedom. Table VI investigates where to apply the consistency loss ($\mathcal{L}_{\mathrm{con}}$) from [17]: either directly in the output space ($\mathbf{Z}$, row 2) or within the camera-conditioned features at each scale ($\mathbf{D}|\mathbf{E}$, row 1). The results indicate minimal differences from applying the loss directly in the output space. Therefore, based on Occam’s razor, we adopt the simpler and more effective design from row 2 as the final approach. Conditioning Impact.  As previously explored in [17], we analyze the impact of our proposed camera conditioning in Table VII. This ablation includes both UniDepth and UniDepthV2 under the same conditions—without $\mathcal{L}_{\mathrm{EG-SSI}}$ and without invariance applied to deep features ($\mathbf{D}|\mathbf{E}$). The results show that conditioning has a even stronger positive effect for UniDepthV2, as evidenced by comparing row 3 vs. row 4 against the comparison of row 1 vs. row 2. Confidence.  The confidence measure introduced in Sec. III-E is evaluated on three zero-shot datasets, as shown in Fig. 5. The y-axis represents the normalized $\mathrm{RMSE}$, computed as $\mathrm{RMSE}$ divided by its per-dataset value at $x=0$, while the x-axis corresponds to the confidence quantile. For each quantile, the evaluation considers only pixels whose confidence exceeds the given threshold. Ideally, confidence should be negatively correlated with error: if the confidence estimate is reliable, higher-confidence regions should exhibit lower $\mathrm{RMSE}$. More specifically, Fig. 5 validates how the predicted confidence of UniDepthV2 negatively correlates with the error, thus showing its reliability.

### v2

@@ IV-B Comparison with The State of The Art
TABLE III: Results for Challenging Domains. All methods are tested in a zero-shot fashion. Missing values (-) indicate the model’s inability to produce the respective output. †: Requires ground-truth (GT) camera for 3D reconstruction. ‡: Requires GT camera for 2D depth map inference.
Method HAMMER Booster FLSea
$\mathrm{\delta_{1}}\uparrow$ $\mathrm{ARel}\downarrow$ $\mathrm{F_{A}}\uparrow$ $\mathrm{\rho_{A}}\uparrow$ $\mathrm{\delta_{1}}\uparrow$ $\mathrm{ARel}\downarrow$ $\mathrm{F_{A}}\uparrow$ $\mathrm{\rho_{A}}\uparrow$ $\mathrm{\delta_{1}}\uparrow$ $\mathrm{ARel}\downarrow$ $\mathrm{F_{A}}\uparrow$ $\mathrm{\rho_{A}}\uparrow$
Metric3D†‡ [14] $0.9$ $122.5$ - - 1.2 $201.2$ - - 0.7 $117.9$ - -
Metric3Dv2†‡ [16] $\underline{65.3}$ $\underline{21.4}$ - - $9.1$ $69.0$ - - $17.5$ $85.7$ - -
ZoeDepth [39] $0.9$ $90.2$ - - $28.5$ $52.7$ - - $15.7$ $161.1$ - -
UniDepth [17] $1.8$ $56.9$ $52.1$ $55.3$ $42.9$ $34.7$ $42.7$ $\underline{73.7}$ $29.9$ $70.2$ $53.6$ $71.0$
MASt3R [51] $2.2$ $69.7$ $38.1$ $\mathbf{86.5}$ $37.9$ $46.1$ $43.9$ $69.6$ $26.7$ $41.5$ $53.6$ $67.5$
DepthPro [43] $29.4$ $37.5$ $\underline{71.0}$ $69.1$ $58.4$ $25.7$ $11.1$ - $10.6$ $145.2$ $4.0$ $4.7$
UniDepthV2-Small $30.6$ $32.1$ $57.0$ $65.6$ $54.4$ $30.3$ $48.6$ $70.1$ $72.3$ $20.6$ $90.5$ $87.2$
UniDepthV2-Base $\mathbf{89.7}$ $22.4$ $68.5$ $\underline{76.5}$ $\underline{62.1}$ $\underline{24.4}$ $\underline{55.1}$ $\mathbf{75.2}$ $\underline{89.1}$ $\underline{12.2}$ $\underline{95.3}$ $\mathbf{93.1}$
UniDepthV2-Large $\underline{64.5}$ $\mathbf{17.5}$ $\mathbf{74.9}$ $\underline{78.3}$ $\mathbf{67.6}$ $\mathbf{19.6}$ $\mathbf{57.9}$ $66.6$ $\mathbf{90.5}$ $\mathbf{12.1}$ $\mathbf{95.9}$ $\underline{92.8}$
@@
Figure 4: Insensitivity to image shape. UniDepthV2 is trained with a variable input shape pipeline in addition to random resizing for each of the image pairs. The proposed training strategy improves the robustness in terms of predicted depth scale and accuracy ($\delta_{1}$) to the input image’s shape compared to two other state-of-the-art methods.
@@
We evaluate our method on a total of ten zero-shot validation sets. The domains cover indoor, outdoor, and challenging scenes, e.g. underwater or transparent objects, shown in Tables I, II and III, respectively. Our model performs better than or at least on par with all baselines, even outperforming methods that require ground-truth camera parameters at inference time, such as [14, 16]. Notably, UniDepthV2 excels in 3D estimation, as reflected in the $\mathrm{F_{A}}$ metric, where it achieves a consistent improvement ranging from 0.5% to 18.1% over the second-best method. Additionally, it outperforms UniDepth [17] in nearly all cases, except for the $\mathrm{\rho_{A}}$ metric on IBims-1, DDAD, and NuScenes. Moreover, we evaluate and compare inference efficiency in Table VIII. All models are run with the same settings. UniDepthV2 is among the fastest and most efficient, even though it requires an additional module for camera prediction. This demonstrates that our proposed version is a significant step forward in both performance and efficiency. However, the camera parameter estimation ($\mathrm{\rho_{A}}$) sees only marginal improvements, indicating that the limited diversity of training cameras remains a challenge that could be addressed with additional camera-only training, as suggested in [43].
TABLE IV: Edge evaluation. Scale-invariant boundary $\mathrm{F_{1}}$ on ETH3D, IBims-1, and Sintel following the protocol of [43]. †: inference performed at fixed resolution of $1536\times 1536$ instead of the input original.
Method ETH3D IBims-1 Sintel
Metric3Dv2 [16] $2.86$ $13.48$ $23.03$
ZoeDepth [39] $0.75$ $4.16$ $2.63$
UniDepth [17] $0.22$ $2.64$ $0.42$
MASt3R [51] $0.60$ $1.45$ $1.08$
DepthPro [43] $\mathbf{4.04}$ $\mathbf{19.38}$ $\underline{29.24}$
UniDepthV2-Small $2.13$ $10.53$ $20.58$
UniDepthV2-Base $\underline{2.99}$ $12.93$ $28.25$
UniDepthV2-Large $2.95$ $\underline{13.75}$ $\mathbf{33.08}$
Table V and Table VI show results for models fine-tuned on the NYU and KITTI training sets and evaluated on their respective validation splits, following standard protocols. Fine-tuning performance serves as an indicator of a model’s ability to specialize to specific downstream tasks and domains. UniDepthV2 effectively adapts to new domains and outperforms methods with similar capacity that were pre-trained on large, diverse datasets before fine-tuning on NYU or KITTI, such as [39, 16, 29]. This is particularly evident in the outdoor setting (KITTI), as shown in Table VI. As detailed in Section III-E, our training strategy incorporates variable image aspect ratios and resolutions within the same distributed batch. Combined with camera conditioning and invariance learning, this approach enhances the model’s robustness to changes in input image shape. Figure 4 quantifies this effect: the y-axis represents normalized metric accuracy ($\mathrm{\delta}_{1}$ scaled by the method’s maximum value), while the x-axis varies the image shape. The normalization ensures a consistent scale across models. UniDepthV2 is almost invariant to image shape, demonstrating that it can effectively trade off resolution for speed without sacrificing accuracy, as clearly illustrated in Figure 4.
[…]
Figure 5: Impact of resolution on memory and runtime. Frames-per-second (top) and peak memory (bottom) versus input resolution in Mega Pixels (log-scale). Missing points at 32MP are due to OOM.
@@
TABLE V: Comparison on NYU validation set. All models are trained on NYU. The first 4 are trained only on NYU. The last 4 are fine-tuned on NYU.
@@
UniDepthV2 -Large $\underline{98.8}$ $\mathbf{99.8}$ $\mathbf{100}$ $\mathbf{4.68}$ $\mathbf{0.180}$ $\mathbf{0.020}$
TABLE VI: Comparison on KITTI Eigen-split validation set. All models are trained on KITTI Eigen-split training and tested on the corresponding validation split. The first 4 are trained only on KITTI. The last 4 are fine-tuned on KITTI.
@@
UniDepthV2 -Large $\mathbf{98.9}$ $\mathbf{99.8}$ $\underline{99.9}$ $\mathbf{3.73}$ $\mathbf{1.71}$ $\mathbf{0.061}$
@@
TABLE VII: Uncertainty evaluation. “In-domain” uses the validation splits of the training datasets; “Zero-shot” evaluates on unseen datasets. $\mathrm{AUSE}$ [86] is the area under the sparsification error curve w.r.t. $\delta_{1}$ score, and $\mathrm{nAUSE}$ is the normalized version where 0 = oracle, 1 = random. $\rho$ is the Spearman rank correlation between predicted uncertainty and error.
Model In-domain Zero-shot
$\mathrm{AUSE}\downarrow$ $\mathrm{nAUSE}\downarrow$ $\rho\uparrow$ $\mathrm{AUSE}\downarrow$ $\mathrm{nAUSE}\downarrow$ $\rho\uparrow$
UniDepthV2-Small $0.020$ $0.221$ $0.682$ $0.040$ $0.539$ $0.289$
UniDepthV2-Base $0.018$ $0.212$ $0.721$ $0.034$ $0.637$ $0.291$
UniDepthV2-Large $0.017$ $0.199$ $0.744$ $0.032$ $0.645$ $0.299$
TABLE VIII: Efficiency results. Inference efficiency results for all methods. Hardware is a A6000 with mixed precision and 0.5 Megapixel images, all methods use a ViT backbone; ViT-Large for competing methods. †: inference on native $1536\times 1536$ resolution. ‡: ConvNext-L backbone.
Method Latency (ms) Params (M) FLOPS (T) Memory (GiB)
Metric3D [14] $29.6$ $203.2$ $0.90$ $1.71$
Metric3Dv2 [16] $133.6$ $411.9$ $3.47$ $3.50$
ZoeDepth [39] $64.8$ $346.1$ $2.08$ $2.02$
UniDepth [17] $91.0$ $347.0$ $2.02$ $2.81$
MASt3R [51] $357.3$ $688.6$ $3.19$ $4.94$
DepthPro [43] $270.6$ $952.0$ $19.3$ $8.42$
UniDepthV2-Small $23.0$ $34.18$ $0.29$ $0.66$
UniDepthV2-Base $35.1$ $114.4$ $0.82$ $1.32$
UniDepthV2-Large $65.4$ $353.8$ $2.17$ $3.47$
IV-C Ablation Studies
The importance of each new component introduced in UniDepthV2 in Sec. III is evaluated by ablating the method in Tables IX, X, XI, and XII. All ablations exploit the predicted camera representation, if not stated otherwise. Table IX evaluates the impact of various architectural modifications compared to UniDepth [17], analyzing their effects on both performance and efficiency. Table X assesses the importance of the proposed loss function (Sec. III-D) and examines the effect of applying the geometric invariance loss originally introduced in UniDepth [17] (Sec. III-C) in different spaces. The rationale behind our design choices is to maintain simplicity while maximizing effectiveness. Additionally, in Table XI we analyze the role of camera conditioning and report results for the original UniDepth under the same training and evaluation setup as our method for a direct comparison. The evaluation is based on four key metrics: $\mathrm{\delta}_{1}$, which measures metric depth accuracy; $\mathrm{SI}_{\log}$, which assesses scale-invariant scene geometry; $\mathrm{F_{A}}$, which captures the 3D estimation capability; and $\mathrm{\rho_{A}}$, which evaluates monocular camera parameter estimation. All reported metrics correspond to the aggregated zero-shot performance across datasets, as detailed in Sec. IV-A.
TABLE IX: Architectural ablations. The different architectural additions (“+”) and subtractions (“-”) from the original UniDepth [17] are reported. “- SHE + Sine”: camera encoding via Sine encoding instead of Spherical Harmonic Transform of the pinhole-based pencil of rays. “- Attention”: attention layers in the decoder are removed. “+ ResNet Blocks”: the attention layers in the decoder are substituted with simpler ResNet blocks. “+ Multi-resol.”: the decoder has lateral connections with the shallower encoder layer, rather than a simpler merging of all resolutions in the bottleneck.
@@ IV-C Ablation Studies
5 + Multi-resol $54.5$ $16.3$ $56.0$ $77.9$ 25.0 34.2
6 UniDepthV2 $60.0$ $15.3$ $57.9$ $79.8$ 25.0 34.2
TABLE X: Loss ablations. $\mathcal{L}_{\mathrm{EG-SSI}}$ refers to either employing or not the proposed Edge-Guided Normalized loss; $\mathbf{O}_{\mathcal{L}_{\mathrm{con}}}$ indicates the output there the geometry consistency loss is applied to.
@@
TABLE XI: Model ablations. The “Model” column refers to architecture and training strategy employed. “V1” is the original UniDepth, while “V2” is the proposed UniDepthV2. “Cond” specifies whether the camera-prompting mechanism is present or not.
@@
TABLE XII: Representation. The “Repr” column refers to direct Cartesian regression (xyz) regression vs. UniDepth’s pseudo-spherical output (sph). The “Model” column refers to architecture and training strategy employed: “V1” is the original UniDepth, while “V2” is the proposed UniDepthV2.
Model Repr Zero-shot Test
1 V1 xyz $48.2$ $21.2$ $41.2$ $61.3$
2 V1 sph $54.5$ $16.4$ $56.1$ $77.1$
3 V2 xyz $49.1$ $20.4$ $38.8$ $60.2$
4 V2 sph $60.0$ $15.3$ $57.9$ $79.8$
Architecture.  Table IX outlines the key modifications that transform the original UniDepth [17] architecture into UniDepthV2. The first major change is the removal of spherical harmonics (SH)-based encoding, which is computationally inefficient. Instead, we revert to standard Sine encoding (row 2). While the difference in performance is minimal in our setup, we hypothesize that the encoding’s impact diminishes as the model benefits from larger and more diverse training data across different cameras. Next, we eliminate the attention mechanism in row 3 due to its high computational cost. This removal results in a significant performance drop, e.g. -4.3% for $\mathrm{\delta}_{1}$, but yields a greater than 2x improvement in efficiency. In row 4, we replace the pure MLP-based decoder with ResNet blocks, introducing spatial $3\times 3$ convolutions. This modification enhances performance by leveraging local spatial structure while inducing a minimal impact on efficiency. Finally, row 5 integrates a multi-resolution feature fusion from the encoder to the decoder, following an FPN-style design. This final architecture significantly reduces computational cost while preserving overall performance: the final model (row 5) achieves similar performance to the original UniDepth (row 1) while requiring only one-third of the computation. Row 6 reports the full UniDepthV2 configuration, i.e. the architecture from row 5 with training augmented with the proposed losses.
$\mathcal{L}_{\mathrm{EG-SSI}}$ Loss.  The effectiveness of the proposed $\mathcal{L}_{\mathrm{EG-SSI}}$ loss, detailed in Sec. III-D, is evaluated in row 2 vs. row 3 of Table X. Introducing this loss results in a 4.7% improvement in $\mathrm{\delta}_{1}$ and a 1.8% improvement in $\mathrm{F_{A}}$, demonstrating its contribution to both metric accuracy and 3D estimation. Interestingly, despite $\mathcal{L}_{\mathrm{EG-SSI}}$ not explicitly supervising camera parameter estimation, the $\mathrm{\rho_{A}}$ metric also shows improvement. This suggests that the loss contributes to a less noisy training process, leading to better feature representations in the encoder. A qualitative comparison of the impact of $\mathcal{L}_{\mathrm{EG-SSI}}$ is presented in Fig. 6. The difference between the third and fourth columns highlights the visual impact of the proposed loss, particularly in refining depth discontinuities. Additionally, the comparison between the second and third columns illustrates the combined effect of architectural changes and increased data diversity, showing improved reconstruction of finer details, such as body parts that were previously smoothed or missed.
$\mathcal{L}_{\mathrm{con}}$ Output Space.  UniDepthV2 introduces multiple instances of camera-conditioned depth features $\mathbf{D}|\mathbf{E}$, corresponding to different decoder resolutions, as described in Sec. III-E. This contrasts with the original UniDepth [17], which relied on a single conditioning point. Given this architectural shift, we argue that deep conditioning may not be optimal. Features at different resolutions encode varying levels of abstraction, and enforcing deep conditioning introduces additional design freedom. Table X investigates where to apply the consistency loss ($\mathcal{L}_{\mathrm{con}}$) from [17]: either directly in the output space ($\mathbf{Z}$, row 2) or within the camera-conditioned features at each scale ($\mathbf{D}|\mathbf{E}$, row 1). The results indicate minimal differences from applying the loss directly in the output space. Therefore, based on Occam’s razor, we adopt the simpler and more effective design from row 2 as the final approach.
Conditioning Impact.  As previously explored in [17], we analyze the impact of our proposed camera conditioning in Table XI. This ablation includes both UniDepth and UniDepthV2 under the same conditions—without $\mathcal{L}_{\mathrm{EG-SSI}}$ and without invariance applied to deep features ($\mathbf{D}|\mathbf{E}$). The results show that conditioning has an even stronger positive effect for UniDepthV2, as evidenced by comparing row 3 vs. row 4 against the comparison of row 1 vs. row 2.
Camera Disentanglement.  We re-evaluate the output representation for UniDepthV2 in Tab. XII, mirroring the analysis in [17]. With otherwise matched camera settings, we observe a similar comparative picture in the UniDepthV2 setting: a significant benefit of our pseudo-spherical representation for depth-specific metrics and a substantial benefit for 3D and camera accuracy.
Confidence.  We evaluate in Tab. VII the confidence estimator introduced in Sec. III-E on the validation splits of the training datasets (“In-domain”) and on unseen datasets (“Zero-shot”). In-domain, uncertainty aligns well with error: $\mathrm{nAUSE}$ is low (0.1990.1990.2210.221) and $\rho$ is high (0.680.680.740.74), both improving as model capacity is increased. Under domain shift, the quality drops, i.e. $\mathrm{nAUSE}$ rises to 0.540.540.650.65 and $\rho$ falls to $0.29$, i.e. indicating that ranking is partly preserved but calibration degrades. The opposing trend, namely slightly higher $\rho$ yet worse $\mathrm{nAUSE}$ for larger models, is probably capacity-driven overfitting: bigger models learn sharper, edge-focused uncertainty priors that order local errors correctly while misestimating their magnitude in novel domains. Nevertheless, even zero-shot uncertainty remains informative and far from random, enabling reliability-aware masking and exploiting our resolution/speed trade-offs at inference.
Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
Figure 7: Failure and Edge Cases. UniDepthV2 is run on challenging images collected from the internet. The domains include mirrors, non-pinhole cameras, paintings, and optical illusions, both from cameras and drawings.
IV-D Failure Cases and Limitations
We present in Fig. 7 six images from the internet that pose unusual, challenging conditions for depth models, including mirrors, paintings, optical illusions, and non-pinhole projection. While UniDepthV2 does not exhibit catastrophic failures on these examples, it can resolve certain human-related optical illusions and correctly recognize a painting when its frame and contextual objects are visible (e.g. “Liberty Leading the People”). However, it struggles with non-realistic paintings and distorted geometry (e.g. “Starry Night”), though it still separates foreground from background. Mirrors remain particularly challenging and ambiguous: without sufficient context, the model often interprets them as just a cavity and accommodates strong deformation. In addition, UniDepthV2 struggles to represent non-pinhole projection and is not able to rectify the deformation as the implicit camera representation is pinhole-based as described in Sec. III-B.